OpenStudy (anonymous):

Bob has $2.90 in pennies, nickels and dimes. He has the same number of pennies and nickels. If the number of dimes is 5 more than the number of nickels, how many of each type of oin does he have?

OpenStudy (anonymous):

thanks guys for helping

OpenStudy (anonymous):

imran's got it

OpenStudy (anonymous):

wait souldnt it be like .05n+.1d+0.01p=2.90

OpenStudy (anonymous):

soooo wat next

OpenStudy (anonymous):

.05n+.1d+.001p=2.90 n=p d=n+5 .05n+.1(n+5)+.01(n)=2.90

OpenStudy (anonymous):

.0001? think it's just .001

OpenStudy (anonymous):

n=15

OpenStudy (anonymous):

yess shouldn't it be that i got 15=x which equals 15 nickels and 15 pennies and 20 dimse

OpenStudy (anonymous):

Wait-I mean.01

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

which equates to the problem because i says p=n and d=n+5

OpenStudy (anonymous):

n=15 , p=15, d=20

OpenStudy (anonymous):

how do u get that answer

OpenStudy (anonymous):

.05n+.1(n+5)+.01(n)=2.90 0.16 n+0.5=2.90 0.16n=2.40 n=2.40/0.16

OpenStudy (anonymous):

n=15

OpenStudy (anonymous):

n=p so p=15 as well

OpenStudy (anonymous):

d=n+5 d=20

OpenStudy (anonymous):

thank you